OpenStudy (volleyballlover55):

Convert 11.42424242 … to a rational expression in the form of a over b, where b ≠ 0. 11 and 42 over 9999 11 and 42 over 9999 11 and 42 over 99 11 and 99 over 42

OpenStudy (volleyballlover55):

@ijakez

OpenStudy (anonymous):

an easy example:\[0.12121212...=\frac{12}{99}\]

OpenStudy (volleyballlover55):

ok sohow do we figure out this question?

OpenStudy (anonymous):

depends on your level of sophistication \[x=.121212...\\ 100x=12.121212... 99x=12\\ x=\frac{12}{99}\] is one method

OpenStudy (anonymous):

type there, i mean to write \[x=.121212...\\ 100x=12.121212...\] subtract get \[99x=12\\ x=\frac{12}{99}\]

OpenStudy (anonymous):

once you do one or two you will see instantly that for example \[0.4444...=\frac{4}{9}\\ .121212...=\frac{12}{99}\\ .123123123...=\frac{123}{999}\] and so on

OpenStudy (volleyballlover55):

so.. i got A

OpenStudy (anonymous):

definitely not A

OpenStudy (anonymous):

two numbers are repeating after the decimal put them over 99

OpenStudy (volleyballlover55):

ok

OpenStudy (volleyballlover55):

i got C! :)